OpenStudy (anonymous):

how to solve using addition, -5x+6y=-12, 5/4x-4=2y

OpenStudy (zpupster):

rearrange the 2nd equation by putting the variables on one side can you do that?

OpenStudy (texaschic101):

also you could multiply the 2nd equation by the LCD, which is 4, to get rid of the fraction

OpenStudy (texaschic101):

Multiply the 2nd equation by 4 and put it in standard form : Ax + By = C

OpenStudy (texaschic101):

5/4x - 4 = 2y -->(4) 5x - 16 = 8y now put it in standard form 5x - 16 = 8y (subtract 8y from both sides) 5x - 8y - 16 = 0 (add 16 to both sides) 5x - 8y = 16 now put you equations together... 5x - 8y = 16 -5x + 6y = -12 --------------add once you add them, what do you get ?

OpenStudy (anonymous):

why am I adding 16 to each side...

OpenStudy (texaschic101):

5x - 8y = 16 -5x + 6y = -12 --------------add -2y = 4 y = 4/-2 y = -2 now sub -2 in for y in either of the equations to find x 5x - 8y = 16 5x - 8(-2) = 16 5x + 16 = 16 5x = 16 - 16 5x = 0 x = 0/5 = 0 check... -5x + 6y = -12 -5(0) + 6(-2) = -12 0 - 12 = -12 -12 = -12 (correct) x = 0 and y = -2 any questions ?

OpenStudy (texaschic101):

you add 16 to both sides because we want the variables on one side and everything else on the other side. We want the equation in standard form.

OpenStudy (texaschic101):

anymore questions ?

OpenStudy (anonymous):

thanks all.. much appreciation

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

i was doing it and got to x=0/5 it confused me.. so i was on the right track... just didn't solv for y yet thanks so much

OpenStudy (texaschic101):

you are correct....x = 0/5 which is the same as 0, so x = 0